# Normalizing flows as an enhanced sampling method for atomistic supercooled liquids

### v1

@@ B.2 Probability distribution of potential energy
We further exploit these data and evaluate the density of state $G(E_{\text{pot}})\propto P(E_{\text{pot}},T)\exp(\beta E_{\text{pot}})$. The interest of the density of state is that it is a temperature independent quantity which is only accurately obtained if proper equilibrium sampling of energy fluctuations is performed. As such it has been used as a tool to assess the degree of equilibration [69].
@@ B.3 Radial distribution function
There are also two different ways to calculate the radial distribution function, $g(r),$ in particle systems at thermal equilibrium. The first traditional approach is based on histograms, and measures the density profile around a tagged particle. The second is based on forces, as recently proposed in Ref. [97]. The identity between both methods is based on the assumption that the system is in thermal equilibrium. Therefore, any difference between the two expressions can be taken as the sign that the system is not equilibrated, but this approach has not be tested before.
Figure 11: Relative difference between two expressions of the radial distribution function $g(r)$, calculated using either the traditional histogram method and the force method [97], for various temperatures. The small systematic differences are due to discretisation issues and do not depend on the degree of equilibration.
@@
This result is reminiscent of similar findings for the configurational temperature, which is shown to decay instantaneously to the thermal temperature during equilibration [98]. The relationship between the histogram and the force methods for $g(r)$ corresponds roughly to a space-dependent generalization of the global configurational temperature.
@@ Appendix C Scaling with system size
Compared to $N=43$, the performance of the MD approach are essentially the same, with deviations appearing near $T_{\rm MD}=0.3$ in both quantities. However, we observe that the efficiency of the three enhanced sampling techniques (PT, PA, and NF) is significantly reduced in larger systems, as expected [68, 20]. In detail, we see that PT and PA now have a comparable performance, with a speedup compared to MD that is much less impressive than for $N=43$ particles. This strong decrease in performance for both techniques stems from the complexity of sampling multiple low-energy states in glassy systems.

### v2

@@ A.1 Hyperparameter
A.2 Ablation study for mixing parameter $\alpha$
Figure 9: Effective sample size $R_{\text{eff}}$, as defined in Eq. (15), for different values of $\alpha$ at $T=0.205.$ The result for normalizing flows (NF) is compared to reweighing (RW), see Sec. III.4.
We have introduced in Sec. IV.2 the parameter $\alpha$ in the loss function which interpolates continuously between energy-based training ($\alpha=1$) and maximum-likelihood training ($\alpha=0$). Which is the best choice for $\alpha$?
To answer this question we have performed different training procedures for different values of $\alpha$. We report in Fig. 9 the effective sample size $R_{\text{eff}}$ which we have identified in the main manuscript as an important factor to quantify the performance of the NF. The figure highlights a maximum near $\alpha\approx 0.5$, which is our final choice. For all other choices, the effective sample size is significantly lower.
In particular, Fig. 9 rules out the possibility to perform pure energy-based training ($\alpha=1$) which would avoid the iterative procedure of finding low temperature training configurations described in Sec. IV.2. In fact, we find that the problem with $\alpha=1$ is not mode-collapse as in other studies in the field of computer vision. For example, we have attempted the $\beta-$NF approach in which the entropy term (i.e., the second term in Eq. (11)) is scaled by a factor $\beta>1$ Sun and Bouman (2021); Higgins et al. (2017) and we did not find any improvement. Instead the only solution we found to increase $R_{\text{eff}}$ for $\alpha\neq 0.5$ is early stopping, which hints to some instabilities in the learning. Nevertheless, even with early stopping $R_{\text{eff}}(\alpha)$ never reaches the the value $R_{\text{eff}}(\alpha=0.5).$
In conclusion, this analysis shows that $\alpha=0.5$ is the optimal choice for the mixing parameter.
@@ B.2 Probability distribution of potential energy
We further exploit these data and evaluate the density of state $G(E_{\text{pot}})\propto P(E_{\text{pot}},T)\exp(\beta E_{\text{pot}})$. The interest of the density of state is that it is a temperature independent quantity which is only accurately obtained if proper equilibrium sampling of energy fluctuations is performed. As such it has been used as a tool to assess the degree of equilibration Yamamoto and Kob (2000).
@@ B.3 Radial distribution function
There are also two different ways to calculate the radial distribution function, $g(r),$ in particle systems at thermal equilibrium. The first traditional approach is based on histograms, and measures the density profile around a tagged particle. The second is based on forces, as recently proposed in Ref. Rotenberg (2020). The identity between both methods is based on the assumption that the system is in thermal equilibrium. Therefore, any difference between the two expressions can be taken as the sign that the system is not equilibrated, but this approach has not be tested before.
Figure 13: Relative difference between two expressions of the radial distribution function $g(r)$, calculated using either the traditional histogram method and the force method Rotenberg (2020), for various temperatures. The small systematic differences are due to discretisation issues and do not depend on the degree of equilibration.
@@
This result is reminiscent of similar findings for the configurational temperature, which is shown to decay instantaneously to the thermal temperature during equilibration Mehri et al. (2021). The relationship between the histogram and the force methods for $g(r)$ corresponds roughly to a space-dependent generalization of the global configurational temperature.
@@ Appendix C Scaling with system size
Compared to $N=43$, the performance of the MD approach are essentially the same, with deviations appearing near $T_{\rm MD}=0.3$ in both quantities. However, we observe that the efficiency of the three enhanced sampling techniques (PT, PA, and NF) is significantly reduced in larger systems, as expected Falcioni and Deem (1999); Earl and Deem (2005). In detail, we see that PT and PA now have a comparable performance, with a speedup compared to MD that is much less impressive than for $N=43$ particles. This strong decrease in performance for both techniques stems from the complexity of sampling multiple low-energy states in glassy systems.